OpenStudy (anonymous):

One third of a number is three more than one fourth of the same number. what is the number?

OpenStudy (anonymous):

\[x = 36\]\[4x = 3x +36\]\[x/3 =(x+12) / 4\]\[x/3 = (x/4)+3\]

OpenStudy (anonymous):

Whoops, don't know why thats upside down. But that is working from the bottom.